OpenStudy (anonymous):

can u explain..how to simplify this... -4^3b^3 x 2b^4

OpenStudy (callisto):

-4^3b^3 x 2b^4 = (-4^3)(2) x b^3 x b^4 = -(64x2)b^(3+4) = -128b^7